OpenStudy (anonymous):

A dog runs 9.5 mi in 0.5 h. The distance the dog travels varies directly with the amount of time it runs. How many miles will it run in 1.75 h? mi

OpenStudy (danjs):

Same exact problem as last one i showed you, different numbers.

OpenStudy (anonymous):

can u help:)

OpenStudy (danjs):

(distance) = k * (time) use (time,distance) = (0.5 , 9.5) Solve for k Use that k in the original equation: (distance) = k * (time) now find the distance when the time = 1.75

OpenStudy (anonymous):

@AlexandervonHumboldt2 can u help me please?

OpenStudy (danjs):

which part do you nto understand?

OpenStudy (danjs):

9.5 = k * 0.5 what is k?

OpenStudy (anonymous):

1.75

OpenStudy (danjs):

9.5 = k * 0.5 9.5 equals k times 0.5

OpenStudy (danjs):

k = 9.5/0.5

OpenStudy (anonymous):

19

OpenStudy (danjs):

now put that k into (distance) = k * (time) and let the time be 1.75 hours, Calculate the distance

OpenStudy (danjs):

33.25 miles

OpenStudy (anonymous):

I got that to thxs:)

OpenStudy (danjs):

np